OpenStudy (anonymous):

Solve: x divided by negative 4- (-8) = 12. I know the answer is 80. But don't know if it should be positive or negative

OpenStudy (anonymous):

Negative or positive 80?

OpenStudy (whpalmer4):

if the signs are the same, the division result is positive, otherwise negative. 4/-2 = -2 4/2 = 2 -4/2 = -2 -4/-2 = 2

OpenStudy (whpalmer4):

is your problem possibly \[\frac{x}{-4} - 8 = 12\]If so, Add 8 to both sides \[\frac{x}{-4}-8+8=12+8\]\[\frac{x}{-4}=20\]Multiply both sides by -4\[\frac{x}{-4}*-4=-4*20\]\[x=-80\]

OpenStudy (anonymous):

No, it's actually x/4- (-8) =12. So the answer would be positive then... thank you :)